# FNOPE: Simulation-based inference on function spaces with Fourier Neural Operators

### v1

@@ 1 Introduction
Probabilistic inference of mechanistic parameters in numerical models is a ubiquitous task across many scientific and engineering disciplines. Among methods for Bayesian inference, simulation-based inference (SBI, [1, 2, 3, 4, 5, 6]) has emerged as a powerful approach for performing inference without requiring explicit formulation or evaluation of the likelihood. Instead, SBI only requires a simulator model which can sample from the likelihood. By training a generative model on pairs of parameters and simulation outputs, SBI can directly estimate probability distributions such as the posterior distribution.
@@
However, existing SBI methods are designed to infer a limited number of vector-valued parameters, which strongly limits their use for inferring spatially and/or temporarily varying, function-valued parameters. In these cases, parameters are commonly inferred on fixed discretizations of the domain. Despite some recent advances leveraging generative models to infer higher-dimensional posterior distributions [7, 8, 9, 10, 11], the high-dimensional inference problems that arise from such approaches remain a challenge.
@@
Furthermore, current models need to be retrained for new discretizations of the parameters or the observations. This is particularly challenging in fields like the geosciences, where observations cannot always be made at the same locations. An alternative to using fixed discretizations is to represent the functions using a fixed set of basis functions, where the inference problem becomes inferring the basis function coefficients, as used, e.g., in [12]. However, these approaches require a good selection of basis functions and suffer from a trade-off between choosing sufficiently expressive basis sets, while maintaining a tractable number of parameters to infer.
@@
To overcome these limitations, we require methods that are capable of modeling and inferring function-valued data. Here, we propose to make use of the Fourier Neural Operator (FNO, [13]) architecture, which operates on function-valued data, for performing SBI on function-valued parameters. Neural operators [14, 15, 16] combine operations on global features of function-valued data with local (typically pointwise) operations, thus capturing both global and local structures. In particular, FNOs use Fourier features to model the global structure. For smoothly varying data, the spectral power is concentrated in the lower frequency components of the spectral decomposition. This allows for a compact representation of the global structure of the data, and hence for the inference of function-valued data on high resolution discretizations.
@@
Figure 1: Overview. FNOPE approximates the posterior over function-valued parameters of a mechanistic model conditioned on function-valued observations. We use a FNO architecture with a flow matching objective to efficiently represent the function-valued parameters, allowing for evaluating extremely high dimensional posterior distributions at arbitrary discretizations of the domain.
@@
We present FNOPE (Fig. 1), an inference method for function-valued parameters: It trains FNOs for Posterior Estimation using a flow-matching objective [17, 9]. FNOPE is capable of solving inference problems for spatially and temporally varying parameters, and can generalize to posterior evaluations on non-uniform, previously unseen discretizations of the parameter and observation domains. Furthermore, FNOPE can estimate additional, vector-valued parameters of the simulator. We demonstrate these features on a collection of benchmark tasks, as well as a challenging real-world task from glaciology. We compare the performance of FNOPE to SBI approaches that use fixed-discretization or basis-function representation of the parameters, and show that FNOPE  outperforms these methods, especially for low simulation budgets. Thus, FNOPE enables efficient inference for high dimensional inference problems that were previously challenging or even intractable.
@@ 2.1 Simulation-based inference
SBI is designed to solve stochastic inverse problems: Given a simulator parametrized by $\theta$, a known prior distribution $p(\theta)$ and an observation $x\in\mathbb{R}^{d_{x}}$, the goal is to infer the posterior $p(\theta\mid x)$ for (typically) vector-valued parameters $\theta\in\mathbb{R}^{d_{\theta}}.$ The simulator implicitly defines the model likelihood $p(x\mid\theta)$ by allowing us to sample $x\sim p(x\mid\theta)$. We can construct a training dataset by sampling from the joint $p(\theta)p(x\mid\theta)$ to construct a dataset of simulations $S=\{(\theta_{i},x_{i})\}_{i=1}^{K}$ for a number of simulations $K$. Standard approaches in neural posterior estimation (NPE) approximate the posterior $q^{\phi}(\theta\mid x)$ with a normalizing flow, which is trained by minimizing the negative log-likelihood $-\mathbb{E}_{(\theta,x)\sim S}\log q^{\phi}(\theta\mid x)$ [1, 3]. In contrast to this, flow-matching posterior estimation (FMPE) [9] learns a conditional velocity field $v^{\phi}_{t,x}(\theta_{t})$ to iteratively denoise samples from a base distribution (typically a Gaussian distribution) to the posterior distribution $p(\theta\mid x)$ The velocity $v_{t,x}^{\phi}$ is trained via the flow matching objective
@@ 2.2 Fourier Neural Operators
We use FNOs [13] to efficiently learn the posterior distribution of function-valued parameters. FNOs are a class of neural operators using the Fourier basis as an intermediate representation of functional data to learn mappings between function spaces. We assume to have a bounded domain $D\subset\mathbb{R}^{d}$, on which we define function spaces $\mathcal{A}(D;\mathbb{R}^{d_{a}})$ and $\mathcal{B}(D;\mathbb{R}^{d_{b}}$). The goal of neural operators is to approximate some given operator $\mathcal{G}:\mathcal{A}\to\mathcal{B}$ by a learnable operator $\tilde{\mathcal{G}}^{\phi{}}:\mathcal{A}\to\mathcal{B}$. In practice, the function-valued data is represented as discretizations of sample functions $a_{i}\in\mathcal{A},b_{i}\in\mathcal{B}$ on the domain $D$.
@@ 3 Method
To extend the standard SBI setting to inferring function-valued parameters, we develop FNOPE by extending FMPE with FNOs as backbone (Figs. 1,2). FNOPE takes the function-valued parameters $\theta$ and observations $x$ as input, and estimates the FMPE flow-field $v^{\phi}$ for function-valued parameters using a combination of several FNO blocks. .
@@ 3.1 Function-valued FMPE objective
To learn the velocity field $v^{\phi}$, we adapt the FMPE objective function [9] for the function-valued setting. Given a discretized observation $(x_{o},l^{x}_{o})$, and a desired parameter discretization $l^{\theta}$, we want to sample ${\theta}\sim{p}_{l^{\theta}}({\theta}\mid{x}_{o},l^{x}_{o})$. This is done by first sampling from a base distribution ${\xi}_{1}\sim{p}_{l^{\theta}}({\xi})$ and then learning the velocity field ${v}^{\phi}_{l^{\theta}}(t,{\xi}_{t},{x}_{o},l^{x}_{o})$ for $\xi_{t}$. In the following, we omit the arguments of $v_{l^{\theta}}^{\phi}$ for clarity. The learned velocity field allows us to iteratively denoise ${\xi}_{1}$ into a sample $\xi_{0}$ from the target posterior distribution. Note that the noise distribution is discretized on the same positions $l^{\theta}$ as the parameter $\theta$. Similarly to Eq. 1, the velocity field ${v}^{\phi}_{l^{\theta}}$ is optimized via the loss function
@@
Here, ${p}_{t,l^{\theta}}({\xi}_{t}\mid{\theta}_{t})$ describes a known noising process such that ${\xi}_{1}\mid{\theta}$ is (approximately) drawn from the base distribution ${p}_{l^{\theta}}({\xi})$. Furthermore, ${u}_{t}({\xi}_{t}\mid{\theta})$ is the true vector field of the path defined by ${p}_{t,l^{\theta}}({\xi}_{t}\mid{\theta}_{t})$. We use the rectified flows formulation [18], such that $u_{t}({\xi}_{t}\mid{\theta}_{t})=({\xi}_{t}-{\theta})$.
@@
The noise distribution, ${\xi}_{1}\sim{p}_{l^{\theta}}({\xi})$, is commonly defined to be independent Gaussian white noise, ${\xi}\sim\mathcal{N}(0,I)$. Such distributions give rise to samples with a uniform power spectrum of $\xi$. As FNOs typically operate on the lower frequency modes of their inputs, independent Gaussian white noise would not be a suitable base distribution choice for our application. Instead, we sample noise from a Gaussian Process $\xi\sim\mathcal{GP}(0,k(\cdot,\cdot))$ [19, 20], where $k(\cdot,\cdot)$ is the square exponential kernel with lengthscale $l$. Using Bochner’s theorem [21, 22], the spectral density of samples is
@@
where $d_{\theta}$ is the dimension of the domain of $\theta$ (and therefore the domain of $\xi$). We choose $l$ to be dependent on the highest Fourier mode $M$ used by the FNO. This ensures that the majority of the signal power in the noise samples $\xi_{1}$ is conserved by the FNO block. We use the heuristic $l=\frac{2}{\pi(M/2+1)}$, which in expectation assures that $>99\%$ of the spectral density of samples $\xi$ is in the lower $M$ frequency modes (derivation in Appendix S3.1). The covariance kernel $k$ of the Gaussian Process is scaled to have unit marginal variance and defines the FMPE noise sampling during training via
@@ 3.2 Adapting to non-uniform, unseen discretizations of the domain
To operate on non-uniform discretizations, we adopt the work of Lingsch et al. [23] and use a FNO with a type II non-uniform fast Fourier transform (NUDFT, [24]). The NUDFT allows us to approximate the first $M$ spectral modes of data or parameters discretized on any $N$ points in the domain. We additionally add the positions $l^{\theta}$ and $l^{x}$ to the input of the FNO blocks through multilayer perceptrons (MLPs, omitted in Fig. 2 for clarity) [25, 26].
@@
In addition to this flexible implementation, we provide and evaluate FNOPE (fix), a variant of FNOPE which uses the fast Fourier transform (FFT) in the FNO blocks and can be used for applications which exclusively consider parameters and observations discretized on uniform grids. The FFT computes spectral components in $O(N\log N)$, compared to $O(N\cdot M^{D})$ for the NUDFT, where $D$ is the dimension of the domain. Therefore, FNOPE (fix) can be scaled to yet higher-dimensional domains (details in Appendix S3.3).
@@ 3.3 Inferring additional parameters
$\mathcal{L}_{2}=\mathbb{E}_{t\sim\mathcal{U}[0,1],(l^{\theta},{\theta},l^{x},{x},\eta)\sim S,{\xi}_{t}\sim{p}_{l^{\theta}}({\xi}_{t}\mid{\theta}),z_{t}\sim p(z_{t}\mid\eta)}||{v}^{\phi}-u_{t}||^{2},$ (4)
@@ 4 Experiments
For the linear Gaussian simulator, we can analytically compute the posterior distribution, allowing us to compare the estimated posterior distributions to this ground truth using the Sliced-Wasserstein Distance (SWD, [27]). However, as is common in SBI applications, we do not have access to ground truth posterior for the other simulators. Instead, we use a combination of two metrics to measure the quality of our posteriors: First, we report the mean square error between predictive simulations from the posterior to test observations. We complement this metric with simulation-based calibration (SBC) [28] on the posterior marginal distributions. We quantify posterior calibration using the Error of Diagonal (EoD), measuring the average distance of the calibration curve of the estimated posterior from a perfectly calibrated posterior (Appendix S2). Good performance on both of these metrics is not a sufficient condition to indicate a correctly estimated posterior, but healthy posteriors typically achieve good performance on these metrics. All evaluation metrics are averaged over three runs and we report mean $\pm$ standard error.
@@ 4.1 Baseline methods
We compare FNOPE to three baseline methods: NPE (with normalizing flows) [3] and FMPE [9] on the coefficients of the spectral basis functions of the parameters (NPE/FMPE (spectral) respectively, details in Appendix S4). We also compare to FMPE with a fixed parameter discretization (FMPE (raw)).
For all baseline methods, we use the sbi toolbox [29]. For the SIRD simulator we compare to Simformer [30], a transformer-based amortized inference approach that is also capable of flexible discretization of function-valued parameters. The other baselines cannot be applied in their basic version to this task because they do not support non-uniform discretizations of both parameters and observations.
@@ 4.2 Linear Gaussian
We first show the ability of FNOPE to approximate the true posterior of a linear Gaussian, as commonly done in SBI benchmarks [31]. To illustrate FNOPE’s ability to infer a large number of parameters, we increase the dimensionality to 1000. We also replace the independent Gaussian prior in this task with a Gaussian process to model smoothly-varying function-valued parameters.
FNOPE clearly outperforms all benchmark methods on this problem (Fig. 3). With a training dataset of $10^{3}$ simulations the SWD is close to zero for both FNOPE  and FNOPE (fix). In contrast, both NPE and FMPE based on spectral features need as many as $10^{5}$ simulations to achieve similar performance. Furthermore, this example shows that the data augmentation applied in training FNOPE, results in a small difference between FNOPE and FNOPE (fix).

### v2

@@ 1 Introduction
Probabilistic inference of mechanistic parameters in numerical models is a ubiquitous task across many scientific and engineering disciplines. Among methods for Bayesian inference, simulation-based inference (SBI, [papamakarios2016snpe, lueckmann2017snpeb, greenberg2019npec, papamakarios2019snle, durkan2020nre, radev2020bayesflow]) has emerged as a powerful approach for performing inference without requiring explicit formulation or evaluation of the likelihood. Instead, SBI only requires a simulator model which can sample from the likelihood. By training a generative model on pairs of parameters and simulation outputs, SBI can directly estimate probability distributions such as the posterior distribution.
@@
However, existing SBI methods are designed to infer a limited number of vector-valued parameters, which strongly limits their use for inferring spatially and/or temporarily varying, function-valued parameters. In these cases, parameters are commonly inferred on fixed discretizations of the domain. Despite some recent advances leveraging generative models to infer higher-dimensional posterior distributions [ramesh2022gatsbi, geffner2022npse, wildberger2023fmpe, schmitt2024consistency, linhart2024diffusionposterior], the high-dimensional inference problems that arise from such approaches remain a challenge.
@@
Furthermore, current models need to be retrained for new discretizations of the parameters or the observations. This is particularly challenging in fields like the geosciences, where observations cannot always be made at the same locations. An alternative to using fixed discretizations is to represent the functions using a fixed set of basis functions, where the inference problem becomes inferring the basis function coefficients, as used, e.g., in [hull2024_sbi_hydro]. However, these approaches require a good selection of basis functions and suffer from a trade-off between choosing sufficiently expressive basis sets, while maintaining a tractable number of parameters to infer.
@@
To overcome these limitations, we require methods that are capable of modeling and inferring function-valued data. Here, we propose to make use of the Fourier Neural Operator (FNO, [li2021FNO]) architecture, which operates on function-valued data, for performing SBI on function-valued parameters. Neural operators [Lu2019DeepONet, kovachki2023neuraloperator, kovachki2024operatoranalysis] combine operations on global features of function-valued data with local (typically pointwise) operations, thus capturing both global and local structures. In particular, FNOs use Fourier features to model the global structure. For smoothly varying data, the spectral power is concentrated in the lower frequency components of the spectral decomposition. This allows for a compact representation of the global structure of the data, and hence for the inference of function-valued data on high resolution discretizations.
@@
Figure 1: Overview. FNOPE approximates the posterior over function-valued parameters of a mechanistic model conditioned on function-valued observations. We use a FNO architecture with a flow matching objective to efficiently represent the function-valued parameters, enabling us to estimate extremely high dimensional posterior distributions at arbitrary discretizations of the domain.
@@
We present FNOPE (Fig. 1), an inference method for function-valued parameters: It trains FNOs for Posterior Estimation using a flow-matching objective [lipman2023flow, wildberger2023fmpe]. FNOPE is capable of solving inference problems for spatially and temporally varying parameters, and can generalize to posterior evaluations on non-uniform, previously unseen discretizations of the parameter and observation domains. Furthermore, FNOPE can estimate additional, vector-valued parameters of the simulator. We demonstrate these features on a collection of benchmark tasks, as well as a challenging real-world task from glaciology. We compare the performance of FNOPE to SBI approaches that use fixed-discretization or basis-function representation of the parameters, and show that FNOPE  outperforms these methods, especially for low simulation budgets. Thus, FNOPE enables efficient inference for high dimensional inference problems that were previously challenging or even intractable.
@@ 2.1 Simulation-based inference
SBI is designed to solve stochastic inverse problems: Given a simulator parametrized by $\theta$, a known prior distribution $p(\theta)$ and an observation $x\in\mathbb{R}^{N_{x}}$, the goal is to infer the posterior $p(\theta\mid x)$ for (typically) vector-valued parameters $\theta\in\mathbb{R}^{N_{\theta}}.$ The simulator implicitly defines the model likelihood $p(x\mid\theta)$ by allowing us to sample $x\sim p(x\mid\theta)$. We can construct a training dataset by sampling from the joint $p(\theta)p(x\mid\theta)$ to construct a dataset of simulations $S=\{(\theta_{i},x_{i})\}_{i=1}^{K}$ for a number of simulations $K$. Standard approaches in neural posterior estimation (NPE) approximate the posterior $q^{\phi}(\theta\mid x)$ with a normalizing flow, which is trained by minimizing the negative log-likelihood $\mathbb{E}_{(\theta,x)\sim S}\log q^{\phi}(\theta\mid x)$ [papamakarios2016snpe, greenberg2019npec]. In contrast to this, flow-matching posterior estimation (FMPE) [wildberger2023fmpe] learns a conditional velocity field $v^{\phi}_{t,x}(\theta_{t})$ to iteratively denoise samples from a base distribution (typically a Gaussian distribution) to the posterior distribution $p(\theta\mid x)$ . The velocity $v_{t,x}^{\phi}$ is trained via the flow matching objective
@@ 2.2 Fourier Neural Operators
We use FNOs [li2021FNO] to efficiently learn the posterior distribution of function-valued parameters. FNOs are a class of neural operators using the Fourier basis as an intermediate representation of functional data to learn mappings between function spaces. We assume to have a bounded domain $D\subset\mathbb{R}^{d}$, on which we define function spaces $\mathcal{A}(D;\mathbb{R}^{d_{a}})$ and $\mathcal{B}(D;\mathbb{R}^{d_{b}}$). The goal of neural operators is to approximate some given operator $\mathcal{G}:\mathcal{A}\to\mathcal{B}$ by a learnable operator $\tilde{\mathcal{G}}^{\phi{}}:\mathcal{A}\to\mathcal{B}$. In practice, the function-valued data is represented as discretizations of sample functions $a_{i}\in\mathcal{A},b_{i}\in\mathcal{B}$ on the domain $D$.
@@ 3 Method
To extend the standard SBI setting to inferring function-valued parameters, we develop FNOPE by extending FMPE with FNOs as backbone (Figs. 1,2). FNOPE takes the function-valued parameters $\theta$ and observations $x$ as input, and estimates the FMPE flow-field $v^{\phi}$ for function-valued parameters using a combination of several FNO blocks.
@@ 3.1 Function-valued FMPE objective
To learn the velocity field $v^{\phi}$, we adapt the FMPE objective function [wildberger2023fmpe] for the function-valued setting. Given a discretized observation $(x_{o},l^{x}_{o})$, and a desired parameter discretization $l^{\theta}$, we want to sample ${\theta}\sim{p}_{l^{\theta}}({\theta}\mid{x}_{o},l^{x}_{o})$. This is done by first sampling from a base distribution ${\xi}_{1}\sim{p}_{l^{\theta}}({\xi})$ and then learning the velocity field ${v}^{\phi}_{l^{\theta}}(t,{\xi}_{t},{x}_{o},l^{x}_{o})$ for $\xi_{t}$. In the following, we omit the arguments of $v_{l^{\theta}}^{\phi}$ for clarity. The learned velocity field allows us to iteratively denoise ${\xi}_{1}$ into a sample $\xi_{0}$ from the target posterior distribution. Note that the noise distribution is discretized on the same positions $l^{\theta}$ as the parameter $\theta$. Similarly to Eq. 1, the velocity field ${v}^{\phi}_{l^{\theta}}$ is optimized via the loss function
@@
Here, ${p}_{t,l^{\theta}}({\xi}_{t}\mid{\theta}_{t})$ describes a known noising process such that ${\xi}_{1}\mid{\theta}$ is (approximately) drawn from the base distribution ${p}_{l^{\theta}}({\xi})$. Furthermore, ${u}_{t}({\xi}_{t}\mid{\theta})$ is the true vector field of the path defined by ${p}_{t,l^{\theta}}({\xi}_{t}\mid{\theta}_{t})$. We use the rectified flows formulation [liu2023rectified], such that $u_{t}({\xi}_{t}\mid{\theta}_{t})=({\xi}_{t}-{\theta})$.
@@
The noise distribution, ${\xi}_{1}\sim{p}_{l^{\theta}}({\xi})$, is commonly defined to be independent Gaussian white noise, ${\xi}\sim\mathcal{N}(0,I)$. Such distributions give rise to samples with a uniform power spectrum of $\xi$. As FNOs typically operate on the lower frequency modes of their inputs, independent Gaussian white noise would not be a suitable base distribution choice for our application. Instead, we sample noise from a Gaussian Process $\xi\sim\mathcal{GP}(0,k(\cdot,\cdot))$ [bilos23_temporal_diffusion, tong2024improving], where $k(\cdot,\cdot)$ is the square exponential kernel with lengthscale $l$. Using Bochner’s theorem (Appendix S3.1) [stein2012gps, williams2006gps], the spectral density of samples is
@@
where $d_{\theta}$ is the domain dimension of $\theta$ (and therefore of $\xi$). We choose $l$ to be dependent on the highest Fourier mode $M$ used by the FNO. This ensures that the majority of the signal power in the noise samples $\xi_{1}$ is conserved by the FNO block. We use the heuristic $l=\frac{2}{\pi(M/2+1)}$, which in expectation assures that $99\%$ of the spectral density of samples $\xi$ is in the lower $M$ frequency modes (derivation in Appendix S3.2). The covariance kernel $k$ of the Gaussian Process is scaled to have unit marginal variance and defines the FMPE noise sampling during training via
@@ 3.2 Adapting to non-uniform, unseen discretizations of the domain
To operate on non-uniform discretizations, we adopt the work of Lingsch2024DSE and use a FNO with a type II non-uniform fast Fourier transform (NUDFT, [Greengard2004NUDFT]). The NUDFT allows us to approximate the first $M$ spectral modes of data or parameters discretized on any $N$ points in the domain. We additionally add the positions $l^{\theta}$ and $l^{x}$ to the input of the FNO blocks through multilayer perceptrons (MLPs, omitted in Fig. 2 for clarity) [Bonev2023SphericalFN, li2023geometryinformed].
@@
In addition to this flexible implementation, we provide and evaluate FNOPE (fix), a variant of FNOPE which uses the fast Fourier transform (FFT) in the FNO blocks and can be used for applications which exclusively consider parameters and observations discretized on uniform grids. The FFT computes spectral components in $O(N\log N)$ in the number of points in the discretization $N$, compared to $O(N\cdot M^{D})$ for the NUDFT, where $D$ is the domain dimension. Therefore, FNOPE scales favorably with the parameter dimensionality $N$ (the number of points in the discretization), but FNOPE (fix) scales favorably with the domain dimensionality $D$ (details in Appendix S3.4).
@@ 4 Experiments
For the linear Gaussian simulator, we can analytically compute the posterior distribution, allowing us to compare the estimated posterior distributions to this ground truth using the Sliced-Wasserstein Distance (SWD) [bonneel2015slicedwasserstein]. However, as is common in SBI applications, we do not have access to the ground truth posterior for the other simulators. Instead, we use a combination of two metrics to measure the quality of our posteriors: First, we report the predictive mean square error (Pred. MSE) between ground truth observations and predictive simulations from posteriors conditioned on those observations. We complement this metric with simulation-based calibration (SBC) [talts2018validating] on the posterior marginal distributions. We quantify posterior calibration using the Error of Diagonal (EoD), measuring the average distance of the calibration curve of the estimated posterior from a perfectly calibrated posterior. Good performance on both of these metrics is not a sufficient condition to indicate a correctly estimated posterior, but healthy posteriors typically achieve good performance on these metrics. All evaluation metrics are averaged over three runs and we report mean $\pm$ standard error. We provide more details on all evaluation metrics in Appendix S2. We provide an overview of training and sampling times, as well as network sizes and computational resources used for all tasks, in Appendix S1.
@@ 4.1 Baseline methods
We compare FNOPE to three baseline methods: NPE (with normalizing flows) [greenberg2019npec] and FMPE [wildberger2023fmpe] on the coefficients of the spectral basis functions of the parameters (NPE/FMPE (spectral) respectively, details in Appendix S4). We also compare to FMPE with a fixed parameter discretization (FMPE (raw)).
For all baseline methods, we use the sbi toolbox [BoeltsDeistler_sbi_2025]. For the SIRD simulator we compare to Simformer [glockler2024simformer], a transformer-based amortized inference approach that is also capable of flexible discretization of function-valued parameters. The other baselines cannot be applied in their basic version to this task because they do not support non-uniform discretizations of both parameters and observations.
@@ 4.2 Linear Gaussian
We first show the ability of FNOPE to approximate the true posterior of a linear Gaussian, as commonly done in SBI benchmarks [lueckmann2021benchmarking]. To illustrate FNOPE’s ability to infer a large number of parameters, we increase the dimensionality to 1000. We also replace the independent Gaussian prior in this task with a Gaussian process to model smoothly-varying function-valued parameters.
FNOPE clearly outperforms all benchmark methods on this problem (Fig. 3). With a training dataset of $10^{2}$ simulations the SWD is close to zero for both FNOPE  and FNOPE (fix). In contrast, both NPE and FMPE based on spectral features need as many as $10^{5}$ simulations to achieve similar performance. Furthermore, this example shows that the data augmentation applied in training FNOPE, results in a small difference between FNOPE and FNOPE (fix), which is an effect of the introduced positional noise. FNOPE learns a posterior under a slightly broader likelihood than what is defined by the model and for very constrained posteriors the posterior quality is slightly poorer. However, we will see that for more challenging tasks, this is an acceptable trade off, as we gain flexibility on evaluation points.
We perform ablation experiments (Appendix S7.1), and observe that the performance of FNOPE is dependent on using sufficiently many Fourier modes in the FNO blocks (Fig. S1a,b). However, other hyperparameters show less influence on the performance (Fig. S1c-e).